### v2

# Labelling Data with Unknown References

Abstract
An evaluator is trustworthy when there exists some agreed-upon way to measure its performance as a labeller. The two ways to establish trustworthiness are either by testing it, or by assuming the evaluator ‘knows’ somehow the way to label the corpus. However, if labelled references (e.g., a development set) are unavailable, neither of these approaches work: the former requires the data, and the latter is an assumption, not evidence. To address this, we introduce an algorithm (the ‘No-Data Algorithm’) by which to establish trust in an evaluator without any existing references. Our algorithm works by successively posing challenges to said evaluator. We show that this is sufficient to establish trustworthiness w.h.p., in such a way that when the evaluator actually knows the way to label the corpus, the No-Data Algorithm accepts its output; and, conversely, flags untrustworthy evaluators when these are unable to prove it. We present formal proofs of correctness and limited experiments.
1 Introduction
One of the most fundamental problems in machine learning, and AI broadly, is measurement. It has been known for a while that some metrics have a certain level of unreliability; for example, by not corresponding to human judgements appropriately [34, 11, 3, 13]. However, the rise of LLMs as evaluators, typically known as LLMs-as-judges, has led to a substantial divide in the field regarding their trustworthiness as a measurement tool. Their use has also led to broader questions around measurement in AI; specifically, about the validity of some results [11, 33, 25].
A key component in the ‘measurement in AI question’, and the object of study of our work, is whether the tool providing the evaluation (the evaluator) may be trusted. This has become more critical given the fast adoption of foundation models in industry and scientific research, and the continuous need to scale. In these cases, trust is usually established based on some sort of ground truth: an existing set of labelled references by which to perform statistical analyses (e.g., a $t$-test). However, it is often the case that said references do not necessarily exist: perhaps the problem is too niche, or the annotation too costly. It follows then that the only way one can establish trust on an unknown evaluator in a no-reference scenario is via formal methods (i.e., formal proofs).
In this paper we introduce an algorithm (which we call the ‘No-Data Algorithm’) by which to establish trust in an evaluator without previously-labelled data. The algorithm is designed to work in-line with contemporary machine learning techniques, where an evaluator claims that it knows the map between datapoints and labels. It is up to a verifier to decide, based on the outcome of certain challenges, whether the evaluator may be trusted. The No-Data Algorithm requires a linear number of calls and does not require labelled datasets (e.g., development or training sets) or knowledge of the labels in the test set to establish trust.
We show that our algorithm, after $r$ calls, bounds confidence on the labeller to $(1/4)^{r}$. This success rate reflects the ‘true’ accuracy of the evaluator: if the evaluator knows how to label the set, the accuracy and success rates will be high; if the evaluator lies about knowing the labels, the accuracy could be high, but the success rate won’t. We present small experiments in synthetic data for traditional machine-learning (i.e., a decision tree) and contemporary (r. an LLM) methods which adjust to the theory’s predictions, although we note our contribution is theoretical in nature. [^1] Code and data may be found in https://github.com/adewynter/no_data_algorithm. Amongst our experimental findings we note that the behaviour of a reliable (read: trained) evaluator adjusts to our theoretical predictions, while an LLM requires more work due to prompting and spurious reasoning. Nonetheless, when properly tuned, it adjusts to the predictions. Our results demonstrate that trust in an evaluator may be established via formal methods without the need for references. A side effect of our experimentation opens a wide range of questions around faithful generation over arbitrary strings for a generative model.
2 Related Work
There has been considerable work on attempting to create reliable evaluators. Most of contemporary work relies on LLMs and various prompting strategies or call stacks, although the problem of (un)reliability/trustworthiness of an evaluator is older than generative models. For example, it has been known that the use of BLEU for tasks other than English-based machine translation is not accurate [22, 28]; and even then, that it does not correspond well to human judgements [34, 11].
More generally, it is known that metric choice, and hence the choice of evaluator, is system-dependent [4, 25, 10, 39, 32]. Poorly-developed rubrics also lead to unreliable human judgements [7, 20], and thus the evaluator trustworthiness problem is also a rubric problem. There are algorithms [27, 47] and paradigms [26] for learning with noisy teachers. Likewise, provable methods have been used with success in other areas, such as representation learning [15], and to improve robustness via synthetic data generation [8, 9]. However, they all assume that a (perhaps noisily) labelled dataset exists.
Relevant to our work, the use of LLMs as evaluators has promised extraordinary scaling capabilities, such as fast data generation and evaluation without humans [29, 6, 5, 23, 41, 48]. Some of these criticisms are that they do not correlate with human judgements well [3, 13, 4, 45, 37]; are sensitive to their prompt [24, 14, 46]; their performance could be the product of memorisation [19, 44, 35]; and even that they do not understand the task at all [43, 40, 46]. Their reasoning capabilities have been put into question, from the results being strongly dependent on the choice of metrics [36, 4], to findings that their output reasoning steps contain spurious reasoning [38, 18, 42]. Most of the work on improving trust, however, relies on prompting strategies or call stacks to other LLMs [21].
The main subroutine from the No-Data Algorithm, which we call the Evaluator-Verifier (EV) protocol is based off the well-known Arthur-Merlin (AM) protocol from Babai [2]. It is a type of zero-knowledge proof; see [12] and [1] for primers on the subject. That said, the EV protocol is different in many regards to the AM protocol, since it is designed to fit arbitrary inputs and the larger No-Data Algorithm. In particular, the EV protocol always reveals its secrets, which sets it apart from applications in theoretical computer science. Likewise, our use of the rubric and the aggregator functions are generalisations to partial functions that are not present (or needed) in previous works.
3 Background
4 Notation
Our work relates to datapoints $x\in X$ and labels $y\in Y$. Whenever the label set $Y$ is unspecified, it is the binary label set $Y=\{0,1\}$. We also let $X\subset\{0,1\}^{n}$; that is, every datapoint is a $n$-bit binary string modelling a phenomenon. The assumptions on $X$ and $Y$ are to simplify our proofs; see Section 9 for a discussion on their relevancy.
Likewise, for a binary string $x\in\{0,1\}^{n}$, we denote its relevant subset set as $S_{x}\subset 2^{x}$. The construction of this specific set will depend on the setup, and will be discussed in detail in Section 6.1.
For two functions $f\colon B\rightarrow C,g\colon A\rightarrow B$ we write the composition $f\circ g\colon A\rightarrow C$ as $fg$.
For two sets $A$, $B$, we denote equality up to isomorphism as $A\cong B$, where $A\cong B$ if $f(A)=f(B)$ but $A\neq B$ for some $f\colon U\rightarrow V$. We also consider the special case of equality up to permutation over subsets as $A\equiv B$, where $A\equiv B$ if $\forall a\in S_{A},\exists b\in S_{B}\,.\,g(a)=g(b)$, for some fixed function $g\colon U\rightarrow V$. We assume all functions to be deterministic throughout, unless stated otherwise.
5 Definitions and Problem Setup
In line with standard learning theory, we assume that there exists an unknown map $f\colon X\rightarrow Y$ that takes datapoints to labels faithfully–that is, $f$ provides the true labelling. Our task is to determine whether a function $E\colon X\rightarrow Y$ (a labeller, or evaluator) is equivalent to $f$, without any knowledge of the map itself.
For that, we introduce four definitions:
A criterion $c\colon X\rightarrow\{0,1\}$ is a map that takes in a datapoint $X$ and maps it to a single bit.
The rubric $C\colon\times_{i,\dots,n}\{0,1\}\rightarrow\{0,1\}^{n}$ takes in the binary outputs from an ordered set of criteria $\mathcal{C}=\{c_{1},\dots,c_{n}\}$ and returns a bitstring that is the concatenation of these outputs: $C=c_{1}|c_{2}|\dots|c_{n}$. We say a rubric is total if it explicitly decomposes nonlinear criteria with arity $>1$ (e.g. the xor $c_{i}=c_{a}\oplus c_{b}$), and denote it as $\bar{\mathcal{C}}$. When the evaluation of said criteria can be done separately (e.g., by testing first $c_{a}$, then $c_{b}$, and then their xor $c_{a}\oplus c_{b}$), we say it is a total evaluation.
The aggregator $\sigma\colon\{0,1\}^{n}\rightarrow Y$ is a function that takes in a binary string of length $n$ (the output from the rubric) and maps it to the label set.
Finally, for two $x,x^{\prime}\in X$, we say that $x$ is similar to $x^{\prime}$ if both $x\cong x^{\prime}$ and $x\equiv x^{\prime}$.
Intuitively, the rubric is the set of reasons why a datapoint $x$ would take a given label (for example, an annotation rubric passed in to human labellers, the last layer in a model, etc.); while the aggregator decides the final label based on that rubric (e.g., via an activation function, a majority vote, etc.).
We are now ready to introduce the core assumption of our work:
Assumption 1.
The map $f\colon X\rightarrow Y$ is a composition of the rubric and the aggregator. That is,
$f=\sigma C.$ (1)
While not a particular strong assumption (after all, datapoints must have a reason to acquire a label and, when there is more than one reason, there must be a way to decide how to aggregate them) this decomposition is crucial to our work.
6 The No-Data Algorithm
The No-Data Algorithm is an algorithm with two components (players): the evaluator and the verifier. [^2] In theoretical computer science, the evaluator is typically known as the prover. We use this more-intuitive naming to highlight the distinction between our work (data-labelling) and others (e.g. theorem-proving). Throughout the algorithm, for every datapoint $x\in X$, the evaluator must convince the verifier that it is able to label $x$ correctly (i.e., that it knows $f$).
For this, the players run a multi-round sub-game which we refer to as the Evaluator-Verifier (EV) protocol. The EV protocol takes in $x$, and either succeeds or ends in failure, w.h.p., based on challenges posed by the verifier on a $x^{\prime}$ similar to $x$ generated by the evaluator. It then returns the proposed label and the status (success or failure).
The algorithm then, in case of failure, decides to flip the label with probability $\phi$. At the end of the loop, it aggregates the results of the EV protocol and returns the final labelling and success count. The final labelling depends on the choice of $\phi$, and the evaluator’s trustworthiness (and hence that of said labelling) is be determined by measuring the success count.
The full algorithm may be found in Algorithm 1. Given that the EV protocol is a fundamental part of the No-Data Algorithm, we focus the remainder of this section to describe it, and provide a characterisation of the No-Data Algorithm (bounds, correctness, etc.) in Section 7.1.
1:  Input: Unlabelled data $X$, evaluator $E$, verifier $V$, criterion $C$, flip probability $\phi$, rounds $r$
2:  $predictions\leftarrow\{\}$
3:  $successes\leftarrow\{\}$
4:  for $x\in X$ do
5:     $success,y\leftarrow\textsc{EV}(x,E,V,C,r)$// Run for $r$ rounds
6:     if $success$ then
7:        $predictions\leftarrow predictions\cup\{y\}$
8:        $successes\leftarrow successes\cup\{1\}$
9:     else
10:        $\tilde{y}\leftarrow\lnot y\text{ with probability }\phi$ // The opposite label
11:        $predictions\leftarrow predictions\cup\{\tilde{y}\}$
12:        $successes\leftarrow successes\cup\{0\}$
13:     end if
14:  end for
15:  return  $predictions$, $successes$
Algorithm 1 The No-Data Algorithm. It works by calling a subroutine (the EV Protocol) to provably determine the trustworthiness of the labels produced by the evaluator. Based on its output (success or failure), the algorithm then determines which label to assign to the corpus, with a stochastic label flip $\phi$ added to it in case of failures.
6.1 The Evaluator-Verifier (EV) Protocol
In the EV protocol both players are given a datapoint $x\in X$, and the rubric $C$, but nothing else. In particular, $\sigma$ is not explicitly given to either player. However, the evaluator claims to know the map $f\colon X\rightarrow Y$. Since $C$ is given, then the claim is that it has prior knowledge of $\sigma$. The implication–were the evaluator to convince the verifier–then is that its choice of label $y$ for $x$ may be trusted.
Remark that the verifier does not know if $y$ is correct. In order to be convinced, the verifier chooses one out of two challenges, neither of which rely on $\sigma$ or $y$. Based on whether the evaluator passes or fails them, the verifier builds its own belief on the evaluator’s trustworthiness.
The entire point of the EV protocol (and thus of the No-Data Algorithm) is for the evaluator to convince the verifier that its choice of label for $x$ is correct. If the evaluator fails, it follows that the correct label for $x$ could be the opposite label. Hence, regardless of success, at the end of the EV protocol the evaluator reveals its label. See Figure 1 for a diagram of the EV protocol. We now describe each of the components and flow.
6.1.1 Setup
At every round of the EV protocol, the evaluator takes in $x$ and generates a similar datapoint $x^{\prime}$ and a ‘partial’ label $\tilde{y}^{\prime}$ that–ideally–is equal to $y$. It is partial because there is no access to $f$, and hence the best the evaluator can do is to generate a guess $\tilde{y}^{\prime}$ based on the encoding $C(x^{\prime})$. If the evaluator indeed knows $f$, an application of $\sigma$ to $C(x^{\prime})$ would lead to the ‘true’ label $y$.
That said, once it has generated the tuple $\langle x^{\prime},\tilde{y}^{\prime}\rangle$, the evaluator may not change them in this round.
Remark that isomorphism is not a sufficient condition for $\tilde{y}^{\prime}=y$ since $C$ might not be total. Hence, both structural and valuation checks are needed. Therefore, after the tuple has been generated, the verifier then chooses uniformly at random one of the two following challenges:
Figure 1: Diagram of the EV protocol. At every round, the evaluator (blue) generates a datapoint $x^{\prime}$ similar to $x$, and a partial label $\tilde{y}^{\prime}$. It then responds to one of two (chosen uniformly at random) challenges by the verifier (orange). The game is repeated by $r$ rounds and returns a succeeded state, unless there is a failure in the challenges, and thus returns a failed state. In either case, it also returns the proposed partial label, $\tilde{y}^{\prime}$.
Challenge 1 (Equality up to Permutation)
Setup: The verifier tests the evaluator by asking it to show a similar point that leads to the same total evaluation of $C$. This challenge is surrogated by testing whether $x\equiv x^{\prime}$ over $\bar{\mathcal{C}}$.
Check: The verifier asserts that
$\forall s\in S_{x}\exists t\in S_{x^{\prime}}\,.\,\forall c\in\bar{\mathcal{C}}\,c(s)=c(t);$ (2)
namely, whether there exists at least one substring in $x^{\prime}$ matching the (total) evaluation of a criterion from $x$.
Challenge 2 (Equality up to Isomorphism)
Setup: The verifier asks for the rationale as to why $f(x^{\prime})=y$. Since $y$ is not available, it checks whether $x^{\prime}\cong x$.
Check: The verifier asserts that $C(x^{\prime})=C(x)$.
6.1.2 Verification
If the evaluator failed the tests, the EV Protocol returns a failure state. Otherwise, it restarts from Section 6.1.1 over and over until the maximum number of rounds $r$ has been reached, or the evaluator fails a test. In either case, at the end of the call the EV Protocol also returns $\tilde{y}^{\prime}$.
Remark 1.
Since the challenges are public, the evaluator could cheat and generate an $x^{\prime}$ to pass one of the challenges, but not both. Otherwise, the implication is that it knows $f$, and it will pass the challenges every time. This is why the challenges assert both structure ($x\equiv x^{\prime}$) and final encoding ($x\cong x^{\prime}$). This observation will be elaborated further in our correctness proofs ( Section 7.1).
Remark 2.
The check for Challenge 1 (demanding equality up to criterion) is very strict. We discuss this in Section 9 and present a relaxed extension in Appendix C.
7 Algorithm Characterisation
In this section we provide proofs of correctness (Section 7.1). Runtime bounds may be found in Appendix B.
7.1 Correctness
Lemma 1 (EV Protocol Correctness Bound).
The probability that the verifier fails to detect a lie by the evaluator in the the EV protocol sub-game after $r$ rounds is $(1/4)^{r}$.
Proof.
(Sketch; a complete proof is in Appendix A.1). We separate the proof in two parts: whether the challenges are necessary and sufficient to determine whether the evaluator is lying; and the probability bound to which it happens.
On the first part, we note that Challenge 1 determines the internal structure of the string, but does not account for the valuation of $C(x^{\prime})$. This challenge is necessary because two $x,x^{\prime}$ could lead to the same evaluation of some $c\in\mathcal{C}$ (e.g., disjunctive clauses), and that would not mean that they are the same up to permutation.
However, Challenge 2 requires strict equality in $C(x^{\prime})=C(x)$, and it is indeed order-dependent, but does not account for said internal structure. Hence both challenges provide some, but not all, information about $x,x^{\prime}$ and the claimed $f(x),f(x^{\prime})$. They are then both sufficient and necessary to determine if an evaluator is lying. More importantly, they are mutually exclusive in the information they provide. This is needed because a lying evaluator could know ahead of time the challenges.
For the second part we handle two cases where the evaluator could lie (or cheat): either (1) it can generate an appropriate $x^{\prime}$, but not a correct $C(x^{\prime})$; or (2) it provides a correct $C(x^{\prime})$ but an incorrect $x^{\prime}$. In either case, the probability that it gets lucky and/or passes the challenge by cheating is $1/4$; after $r$ rounds, it is $(1/4)^{r}$. ∎
Remark 3.
The exponential bound from Lemma 1 means that, for example, after running the EV protocol for 3 rounds, the probability that the verifier fails to detect a lie from the evaluator is 1.6%.
Theorem 1 (No-Data Algorithm Correctness Bound).
Suppose an evaluator has accuracy $\alpha$ on a balanced dataset $X$ of size $n$ with binary labels $Y=\{0,1\}$. Assume the evaluator always lies on any $x\in X$ iff it incorrectly labels the datapoint. Then if the EV protocol is ran for $r$ rounds, the expected accuracy of the No-Data Algorithm is given by
$\mathrm{E}[\text{correct}]\leq 1-(1-\alpha)\left(1-\phi+\phi\left(\frac{1}{4}\right)^{r}\right),$ (3)
for a chosen value of $\phi$ and over the same dataset $X$.
Proof.
(Sketch; a complete proof is in Appendix A.4). The accuracy of any evaluator is directly tied to its success in the algorithm: if it knows the label, then the checks by the EV protocol will pass. Even under the constraint that the mislabelled datapoints are tied to the evaluator’s lies, errors could still get past undetected. From Lemma 1, we know the error bound for every label flip, and therefore the probability of mislabelling the entire dataset. We may compute the latter via an application of the union bound, that is, $\Pr[\text{error}]\leq(1-\alpha)(1-\phi+\phi(1/4)^{r})$. ∎
The implication of Theorem 1 is that a good evaluator with low probability to flip the labels will obtain a good label accuracy. A bad evaluator with high probability to flip the labels will also obtain a good accuracy. Another interpretation of this, however, is that there is no way to recover unknowable labels when the evaluator is unable to convince the verifier–we discuss this further in Section 9.
It is worth mentioning that the assumption for Theorem 1 around always lying when mislabelling (and hence always being wrong) are not necessarily applicable to all evaluators, since a particularly tricky evaluator could still guess the right label and yet not pass the challenges. We consider this out of scope, since the ability to guess a label without explaining how is not a viable way to establish trust in scenarios without any references. Consequentially, for the No-Data Algorithm it is more helpful to look at the success count.
8 Experiments
We describe the two experiments and results in this section and refer the reader to Appendix D for a more detailed treatment, including prompts and specifics on the models used. Throughout this section, we run the No-Data Algorithm with $r=3$.
8.0.1 Data
We created two mostly disjoint synthetic datasets so that we could control the labels, data, and their rubrics. We refer to these datasets as in-phenomenon, or IP; and out-of-phenomenon, or OOP, as opposed to in-distribution (r. out-of-distribution) to highlight the fact that the entire measurement is altered by the change in data–and, that, semantically, we are modelling different phenomena. The IP dataset is our baseline: the evaluator will not be lying on knowing this phenomenon. The OOP dataset is our reference for a corpus that the evaluator has not seen before (and thus it will lie about knowing it). For both problems, $\sigma$ is the majority vote.
We chose to generate our own data to ablate out any potential memorisation concerns, and to control the rubrics. These may be found in Table 1. The test sets for both IP and OOP included 498 entries and are balanced.
8.0.2 Evaluator
We used a decision tree (DT) for our evaluation, and an LLM (o3-mini; [31]). Both models observed a balanced training split of the IP dataset: the DT was trained on it, while the LLM observed five exemplars in the prompt. The rationale to train the DT was to fully ensure that the evaluator learnt both $\sigma$ and $C$. For generation it only had access to the rubric, but not $\sigma$. The LLM observed it via the prompt, and was made aware of the aggregation function only for the labelling step. For generation it was asked to pick the best candidate from the dataset, as opposed to outputting a new point, as we observed marked performance decreases otherwise. During OOP, the LLM was given the rubric for IP and an OOP datapoint; while the DT was trained in IP (and, ideally, learnt the rubric), and classified IP and OOP datapoints.
Rubric Definition
In-Phenomenon
$c_{0}$ $y=1$ if $x$ has an even number of ones; else 00
$c_{1}$ $y=1$ if $x$ starts with a zero or contains $10101$, but not both; else 00
$c_{2}$ $y=1$ if $x$ has strictly more than five ones; else 00
Out-of-Phenomenon
$c_{0}$ $y=1$ if $x$ contains the substring $111$; else 00
$c_{1}$ $y=1$ if $x$ ends with a one; else 00
$c_{2}$ $y=1$ if $x$ contains the substring $110001$; else 00
Table 1: Rubrics and criteria for our experiments. For both problems, $\sigma$ was the majority vote. Remark that $c_{1}$ for the IP rubric is an exclusive-or operation. This means that, while this rubric contains three criteria, the total rubric contains five: $\{c_{0},c_{2},c_{1,a},c_{1,b},c_{1}\}$; where $c_{1,a},c_{1,b}$ are the clauses for $c_{1}$, and $c_{1}$ is the exclusive-or itself.
8.0.3 Verifier
The verifier is a rule-matching algorithm that has access to the relevant rubrics (e.g., when testing IP, it observes the rubric for IP), and works as described in Section 6.1.
8.1 Results
The results are in Table 2. Both the DT and the LLM adjusted to the predictions correctly, successfully convincing the verifier when it was an oracular generator, reaching approximately their knowable accuracy, and outputting low success rates when lying. Indeed, the accuracy of the models in the OOP dataset hovers around 54%–close to random guessing. A similar accuracy can be obtained by the output of the No-Data Algorithm, but the success rates between IP (100%) and OOP (5%) provide better information as to in which situations the evaluator is trustworthy. The tuned $\phi$ value, based on the known labels, also yielded accuracies close to the expected score from the known dataset. Of note, which we will cover in Section 9, is the number of flips performed by the algorithm in the DT OOP dataset: 46%.
As part of our experimentation, we performed ablation studies on the relationship of lying with evaluators and generators (Section 8.1.1); and the effectiveness of generator strategies in LLMs (Section 8.1.2).
8.1.1 Ablation: Lying Evaluators and Lying Generators
The results are in Table 3. Throughout that study DT randomly (w.p. $1/10$) flipped the label when evaluating. The lies tested were (in order in the table) when the model claims to
  1. 1.
    know $\sigma$: that is, it can output an encoding $C(x)$, but does not necessarily know $\bar{C}(x)$;
  2. 2.
    know $f$: that is, it can retrieve a datapoint $x^{\prime}\in X$ such that $y^{\prime}=y$, but nothing else; and
  3. 3.
    approximately know $f$ up to a probability $p=1/10$ (r. it could pass both challenges with probability $1-p$).
The last ‘lie’ is more akin to how a realistic generator would behave, sometimes understanding the challenges well, but failing them with some probability.
As before, the results indicate that lying evaluators tend to score low in the success rate, even when their accuracies are close to the known case. Remark that (a) the evaluator that lies on knowing $\sigma$ had a success rate much higher than that of lying of knowing $f$; and (b) the evaluator with approximately knowledge of $f$ had a comparatively low (44%) success rate. We discuss both findings in Section 9.
DT (known) DT (unknown) LLM (known) LLM (unknown)
IP
Successes/Flips — / — 100.0 / 0.0 — / — 81.3 / 1.8
Accuracy/F1 62.2 / 58.8 62.2 / 59.8 99.8 / 99.8 97.6 / 97.6
OOP
Successes/Flips — / — 4.8 / 46.4 — / — 28.0 / 6.0
Accuracy/F1 54.2 / 54.2 52.8 / 52.1 60.6 / 66.2 59.0 / 64.0
Table 2: Results for the known (test score if the labels were known) and unknown (output of the No-Data Algorithm) cases. The known case is a reference that in practice does not exist. For the unknown case, if we used the base generator, the performance is exactly the same as in the known case. The flip $\phi$ is set near the accuracy in the known case ($0.6$ for DT; $0.9$ for LLM). This is because the generator acts as an oracle, convincing the verifier every time, and hence the No-Data algorithm reduces to simple evaluation (the known case). However, the success rate is extremely low in DT OOP (unknown): even though there is no way for us to know whether the evaluator’s score is trustworthy, the success rate indicates deception.
IP (Flips) IP (No Flips) OOP (Flips) OOP (No Flips)
No Lie: DT
Successes / Flips 100.0 / 0.0 100.0 / —- 6.0 / 35.3 6.4 / —-
Accuracy/ F1 60.6 / 58.8 59.0 / 57.4 48.6 / 47.3 45.6 / 44.8
Lie 1: $\sigma$ unknown
Successes / Flips 17.0 / 35.1 15.1 / —- 2.8 / 35.1 2.2 / —-
Accuracy/ F1 58.6 / 58.8 42.6 / 45.0 52.4 / 53.8 49.0 / 49.4
Lie 2: $f$ unknown
Successes / Flips 0.6 / 37.4 0.2 / —- 3.0 / 35.7 2.8 / —-
Accuracy/ F1 50.6 / 48.1 39.4 / 42.4 51.4 / 52.2 51.2 / 52.8
Lie 3: generator w.p. $p=1/10$
Successes / Flips 43.6 / 24.9 51.3 / —- 4.4 / 37.9 4.6 / —-
Accuracy/ F1 55.8 / 54.2 49.0 / 49.4 47.6 / 49.5 46.6 / 48.6
Table 3: Performance for a lying DT evaluator on the No-Data Algorithm with $\phi=0.6$. From top to bottom: Lie 1: $\sigma$ unknown. Lie 2: $f$ unknown. Lie 3: $f$ approximately known (lying with probability $p=1/10$) in the generation step. In all these calls, the evaluator itself also outputs the incorrect label with probability $p=1/10$. Note the comparatively high success rates in the IP case of Lies 1 and 2. We discuss these findings in Section 9. Also note the consistently low success rate for OOP across all experiments, thus indicating the robustness of this algorithm to any evaluator strategy in scenarios where the data is completely unknown to the evaluator.
8.1.2 Ablation: LLM Generation Strategies
Given the LLMs’ ability to follow instructions, it is likely that their performance could be sensitive to how the datapoints are generated during the EV protocol. Hence in this study we compared the performance of both the LLM whose results we presented in the main results (o3-mini) with another LLM, GPT-4o [30]. We tested two different generation strategies: picking a matching datapoint from the data, as in Section 8.1, and actually generating a datapoint from the rubric. The results are in Table 4. We found that, again, the experimental results adjusted to the predictions, with lying and low-performing models being easily spotted by the No-Data Algorithm. However, we also found that actual generation led to bad performance in both LLMs. We note that this is expected: in general, classifying an element in a language is much easier than generating an element [17, 16].
o3-mini (IP) o3-mini (OOP) GPT-4o (IP) GPT-4o (OOP)
Picking
Accuracy/F1 97.6 / 97.6 59.0 / 70.0 58.0 / 67.1 56.4 / 65.4
Successes/Flips 81.3 / 1.8 27.9 / 6.0 28.7 / 71.3 10.8 / 8.8
Generating
Accuracy/F1 94.2 / 94.2 60.6 / 65.5 56.6 / 63.8 53.4 / 63.6
Successes/Flips 55.4 / 5.0 28.7 / 7.8 5.8 / 10.8 34.5 / 4.6
Table 4: Results of our ablation study on LLM generation strategies, comparing a high-performing model (o3-mini) with a lower-performing model (GPT-4o) as a LLM-based baseline. We report the highest-performing run at $\phi=0.9$. While both models were able to broadly adjust to their original accuracy, their success rates are markedly different when the models are generating a datapoint versus picking one.
9 Discussion
9.1 Theory
Theorem 1 states that when the labeller has high accuracy, a low $\phi$ will be able to reproduce the label set. It is then possible to measure the successes to establish a confidence bound on the aggregate labels provided by the evaluator. However, this result has a different interpretation: when the accuracy of the evaluator is zero, there is no way to determine the labelling, regardless of $\phi$–it is completely unknowable. We argue this interpretation also makes sense: otherwise, it would be possible to create information out of thin air (that is, without any prior knowledge of $f$). Moreover, the primary purpose of the No-Data Algorithm is to establish trustworthiness on the evaluator, not label the data.
One important limitation of the theory is that we assumed that $Y$ is a binary set, and that $X$ is the set of $k$-bit binary strings. The first assumption can be addressed by ‘binarising’ a set $Y$ with arity $>2$ by separating it into two classes. The second, however, must be understood more carefully: if this were a natural-language problem, it would be inaccurate to assume that the $k$-bit encoding for $x\in X$ will retain correctness by solely working with the binary encoding of the strings. Extra information (e.g., semantics captured in a latent space) must be added to $x$, before proceeding as usual.
9.2 Experiments
The results adjusted to the theory’s predictions. In particular, in both the known/unknown comparison and the lying ablation study, the LLM and DT evaluators consistently had low success rates for the OOP case, thus indicating that the No-Data Algorithm is robust to evaluator strategies attempting deception. Three things we noted throughout our experimentation were the high number of flips in the DT OOP case; the (relatively) high success rate when the evaluator lied about knowing $\sigma$; and the low success rate of the evaluator with approximate knowledge of $f$.
The first illustrates our interpretation of Theorem 1: since the model’s accuracy is near-random in the OOP dataset–as indicated by the known case–the verifier will almost always catch a lie. Since the flip probability is set to $\phi=0.5$, half the labels got flipped and no information was gained from the labelling. However, we learnt that the evaluator cannot be trusted–the main point of the No-Data Algorithm.
In the case of the evaluator lying about $\sigma$, the explanation is simple. This success rate depends on the cardinality of $\bar{C}$: when it is equal to $\mathcal{C}$, the model will always pass Challenge 1 (and fail Challenge 2). However, when $\lvert\bar{C}\rvert>\lvert C\rvert$, there will be criteria that the evaluator cannot possibly solve without knowing how $C(x)$ is constructed based on its internal structure, and hence it will fail them with some probability. In our experiments, the internal structure depends on both clauses in $c_{2}$.
Finally, the (relatively) low success rate of the evaluator with approximate knowledge of $f$ illustrates one critical characteristic of the No-Data Algorithm: the success rate is a very sensitive metric, highly dependent on both the evaluator’s labelling and generation reliabilities.
10 Conclusion
We introduced an algorithm to determine trustworthiness of an evaluator in the case where labelled data does not exist. It works by performing checks that we showed to be sufficient and necessary to determine (w.h.p.) whether the evaluator knew the labels. Experimentally, the results adjusted to the theoretical predictions, with the No-Data Algorithm being able to determine lying evaluators under various scenarios. Of note was when the evaluator was randomly guessing the labels: while its accuracy was deceptively close to that of its known case, the success rate of the No-Data Algorithm was low, thus flagging that said labelling could not be trusted.
One natural implication of the No-Data Algorithm’s correctness proof is that, when the evaluator cannot be fully trusted, the data cannot be annotated. We argued that this scenario is unsolvable: there is no prior information on the datapoint, and hence it will be effectively creating information out of thin air. This is why the algorithm’s focus is less on the labelling returned, and more on the successes: if an evaluator has low accuracy, but high success rate, it is very likely it is not lying; just wrong (e.g., the IP case in the experiments). Alternatively, high accuracy but low success rate indicates deception (r. the OOP case). The addition of the flip probability $\phi$ and the success count allows for some–limited–heuristics by which the No-Data Algorithm can produce a labelling when the evaluator’s output is rejected. Nonetheless, since the primary purpose of this algorithm is to establish trustworthiness, and not to label the set, it is up to the user to decide (to its own margin of trust) whether to accept or reject the labelling.
As mentioned, some of our work’s limitations are solvable (e.g., label sets of arity $>$ 2), and others are natural (the interpretation of Theorem 1). However, other limitations are more practical in nature and should be the subject of further work. Namely, it is well-known that LLMs are sensitive to prompting. Establishing trust in an LLM evaluator in our setup was equivalent to establishing trust in the prompt, not the model. This nuance could require extensions to the No-Data Algorithm to determine (w.h.p.) whether it is the model or the prompt what is failing to convince the verifier.
11 Limitations
Most of the limitations in the theory are covered in Section 9, and relevant extensions to more common setups are provided in Appendix C. Even when our focus was not experimentation, it could be argued that these were limited in size and scope. We agree–time constraints made us prioritise rigour over experimental results. We have open-sourced the code and data for our work, but also plan to extend it ourselves with other datasets and open models in the near future.
12 Ethics
We argue that the nature of our work precludes any misuse.
Appendix A Detailed Proofs
In this section we provide more careful treatments of the proofs of Lemma 1 (Appendix A.1) and Theorem 1 (Appendix A.4).
A.1 Correctness of the EV protocol
The key in this proof is to note that accepting a label does not mean it will be the correct label, but that the challenges in the EV protocol are sufficient to ascertain said correctness with some probability.
We separate our proof in two: the first part concerns itself with correctness, or whether the challenges will provide sufficient information to determine if the evaluator can be lying. The second part deals with robustness under lying, cheating, or lucky scenarios by the evaluator.
A.1.1 Part 1 (Correctness)
We prove the following lemma:
Lemma 2.
Let $x,x^{\prime}\in X$ be two binary strings such that $x\neq x^{\prime}$. In order to test whether an evaluator knows $f(x)=f(x^{\prime})$ without calling $f$, it is necessary to challenge the evaluator in both $x\equiv x^{\prime}$ and $x\cong x^{\prime}$. Moreover, these conditions are sufficient.
Proof.
(Necessary) It follows from the definition of $C$. Since $f(x)=f(x^{\prime})$, testing $x\equiv x^{\prime}$ is a necessary condition, but not sufficient. This challenge yields information around the internal structure of $x,x^{\prime}$ up to permutation. It is necessary because it is possible to have two $x,x^{\prime}$ such that $x^{\prime}\not\in\{\pi_{i}(x)\}$ (i.e., two strings that aren’t permutations of one another), and yet $C(x^{\prime})=C(x)$. However, this does not prove that x is similar to x’, since there could be a subset $t\in S_{x^{\prime}}.t\not\in S_{x}$ and $c(t)=c(s)$ for some $c\in\mathcal{C}$ (for example, if $c$ contains a disjunctive clause). Hence running this challenge provides information as to whether the encoding of $C(x^{\prime})$ contains all information fed into $\sigma$ for a specific $x,x^{\prime}$. It accounts for variation in internal structure of the string so that $C(x)=C(x^{\prime})$, but does not account for the ordering of these bitstrings.
On the other hand, $x\equiv x^{\prime}$ will determine whether the encoding for the bitstrings lead to the same valuation, by checking strict equality. This is because, under our setup, the definition of similarity for $x,x^{\prime}$ implies $\sigma C(x^{\prime})=\sigma C(x)$. This makes it order-sensitive, but does not account for the internal structure of the string.
Since neither of them provide information about the other, it follows that both are needed to fully determine $f(x)=f(x^{\prime})$.
(Sufficient) It follows directly from the earlier statement. Suppose there exists a pair $x,x^{\prime}$ such that either (or both) $x\not\equiv x^{\prime}$, $x\not\cong x^{\prime}$ and also $f(x)=f(x^{\prime})$. This would imply that (1) either the string , a contradiction.
This concludes the proof. ∎
While a natural consequence of Lemma 2 is that only checking for isomorphism is required for testing equality, we note that the goal of the challenges is to determine if the evaluator is lying, not the label itself.
A.1.2 Part 2 (Robustness)
We prove the following lemma:
Lemma 3.
The probability of error (i.e., the verifier accepts a $y^{\prime}\neq y$) at the $r^{\text{th}}$ round is $\left(1/4\right)^{r}$.
Proof.
In here we have two scenarios: either the evaluator knows $f$, or it is lying or cheating.
Knowing $f$ implies that the evaluator will generate a ‘good’ $x^{\prime}$ every time, such that $\tilde{y^{\prime}}=y$; and hence it will pass the checks from the verifier w.p.1. For the proof we focus then on the second case: that the evaluator is lying or cheating.
There are then two (disjoint) possibilities for a generated $\langle x^{\prime},y^{\prime}\rangle$: either that the evaluator is lying, but gets lucky; or that it is cheating and can generate the answer for one of the challenges. It is worth noting that generating an answer to pass both challenges is equivalent to knowing $f$.
A.2 Lie 1: Generate a ‘good’ $x^{\prime}$
Suppose that the evaluator generates an $x^{\prime}$ such that $S_{x^{\prime}}=S_{x}$, and thus can pass Challenge 1. It cannot pass Challenge 2, since $S_{x^{\prime}}=S_{x}$ implies that either $C(x^{\prime})\neq C(x)$ (and hence it fails it); or $C(x^{\prime})=C(x)$ (and hence either it gets lucky, or it knows $f$). Since the challenges appear with probability $1/2$, the likelihood that the evaluator gets lucky in this round is $1/4$.
A.3 Lie 2: $f(x^{\prime})\neq y^{\prime}$
Now suppose that the evaluator anticipates Challenge 2, and picks an $x^{\prime}$ such that $C(x^{\prime})=C(x)$. Similar to Lie 1, then either $S_{x^{\prime}}\neq S_{x}$ (and thus it fails Challenge 1); or $S_{x}=S_{x^{\prime}}$ (and hence it gets lucky or it knows $f$). The likelihood of getting lucky this round is again $1/4$.
Therefore, the probability of the verifier to not catch the evaluator’s lie is $1/4$ in any round. At the $r^{\text{th}}$ round, $\left(1/4\right)^{r}$. This concludes the proof. ∎
Therefore, from Lemma 2 we know that both challenges together are sufficient to ascertain the truthfulness of a label, and from Lemma 3 that the ability of the verifier to do this is bounded. Putting both lemmas together we obtain the proof for Lemma 1.
A.4 Correctness Bounds of the No-Data Algorithm
By a simple application of the union bound on the probability that the evaluator was wrong, but not detected by the verifier, and the probability that the evaluator was wrong, detected by the verifier, but without having its label flipped.
For this, it is much easier to work with the errors $\epsilon=1-\alpha$, and remark that, from Lemma 1 the first term (undetected, no flip) is given by
$\Pr[\text{undetected}]$ $=\sum_{f(x)\neq y}\left(\frac{1}{4}\right)^{r}\text{, and hence}$ (4)
$\mathrm{E}[\text{undetected}]$ $=\epsilon\left(\frac{1}{4}\right)^{r}.$ (5)
Likewise, the second term accounts only for the situation where the mislabelling was detected and the label was not flipped:
$\mathrm{E}[\text{detected, no flip}]=\epsilon\left(1-\phi\right)\left(1-\left(\frac{1}{4}\right)^{r}\right).$ (6)
This yields:
[…]
$\leq\left(\frac{1}{4}\right)^{r}+\left(1-\phi\right)\left(1-\left(\frac{1}{4}\right)^{r}\right)$ (8)
$\mathrm{E}[\text{wrong}]$ $\leq\epsilon\left(1-\phi+\phi\left(\frac{1}{4}\right)^{r}\right)\text{, by linearity of expectation.}$. (9)
Substituting back $\Pr[\text{wrong}]=1-\Pr[\text{right}]$ and $\epsilon=1-\alpha$ yields the desired value. This concludes the proof.
Appendix B Characterisation Bounds
Lemma 4.
Suppose the evaluator and verifier run in polynomial time for any input $x\in X$. Then, if the No-Data Algorithm is ran over a dataset $D$; and with $r$ rounds of EV protocol, the runtime is then $\operatorname{O}\!\left(r\lvert D\rvert\right)$.
Proof.
Straightforward. ∎
Appendix C Extensions
C.1 Distance-based Constraints
The constraint that $C(x)=C(x^{\prime})$ is too strict for some applications. We sketch out in this section a variant of the No-Data Algorithm where, instead of requiring equality, we request closeness with respect to some arbitrary metric $\delta$ in Challenge 2. Suppose now that the failure is given whenever $\delta(C(x),C(x^{\prime}))\geq\epsilon$, for some $\epsilon\in[0,1/2]$. The bounds of correctness are roughly equivalent to these described in Lemma 1 and Theorem 1. However, the term $(1/4)^{r}$ now takes a dependency on the probability of generating a string that is far from the desired constraint.
C.2 $k$-ary Label Sets
To convert a dataset of arity $k>2$, it is necessary to rerun the No-Data algorithm $k-1$ times. On each run, the label set $\{l_{1},\dots,l_{k}\}$ is mapped as $\{l_{1}\},\{l_{2},\dots,l_{k}\}\mapsto\{0\},\{1\}$. In these cases, the runtime of the algorithm is increased by a factor of $k-1$.
Appendix D Detailed Experimental Methods
D.1 Detailed Setup
For the LLM we used o3-mini (version 2025-01-31), and called it via the Azure OpenAI API. We set the maximum tokens to 50,000 and left all other parameters as default. For the decision tree we used the implementation from sklearn [^3] https://scikit-learn.org/stable/index.html with the default parameters. All data analysis was performed on a consumer-grade laptop. Random seeds are set for the training in the code.
D.2 Prompts
The prompts we used for our LLM stack are in Prompt 1 (evaluation) and Prompt 2 (generation). All exemplars are omitted for brevity. Whenever a parsing error occurred, we retried the call up to five times. If it failed on the fifth, we returned a random string.
You are labelling binary strings based on a rubric (given below).
First return the parts of the criterion that match the string, and the values.
Then return the label based on the aggregate function.
# Rubric:
<NL RUBRIC GOES HERE>
# Aggregation Function:
Majority vote
Return your answer in the form:
|reasons|
(list of reasons)
|reasons|
|label|
(the label)
|label|
Table 1: System prompt for evaluation. Anchors (‘|reasons|’, ‘|label|’) are used for parsing. The text contained within ‘|reasons|’ is, as specified by the exemplars (not pictured) a chain-of-thought approach determining criterion-by-criterion matching. The rubric, explained in natural language, is inserted in the marker. Throughout the experiments we only used the IP rubric.
You are given five datapoints and a set of criteria.
Based on these, pick the datapoint that matches the criteria exactly.
# Rubric:
<NL RUBRIC GOES HERE>
# Datapoints:
<DATAPOINTS GO HERE>
Return your rationale, and then your chosen datapoint ONLY in the form:
|datapoint|
(the datapoint)
|datapoint|
Table 2: System prompt for generation. This prompt does not utilise exemplars. The rubric, explained in natural language, is inserted in the marker. Throughout the experiments we only used the IP rubric and selected datapoints from the IP training dataset, ensuring that at least one of the entries matched the criterion. Given that the model did not follow the output format exactly, some trial-and-error was needed to ensure a comparatively high parse success rate.
You are a datapoint generator over binary strings.
Given a rubric (given below), a datapoint, and a label, return a similar datapoint that has the same
label, and fulfils the same conditions as the rubric.
For convenience, always start with the same datapoint: 0100. It will be easier to work with.
# Rubric:
<NL RUBRIC GOES HERE>
Return your rationale, and then the final datapoint in the form:
|datapoint|
(the datapoint)
|datapoint|
Table 3: System prompt for generation used in the ablation study from Section 8.1.2. Unlike Prompt 2, this prompt does take exemplars (not pictured): the first five entries in the training dataset, along with chain-of-thought-style rationale.